Give exact and approximate solutions to three decimal places.
y^2-16y+64=36

Answer by mananth(16949)

y^2-16y+64=36
-36
y^2-16y+64-36=0
..
y^2-16y+28
...
y^2-14y-2y+28=0
y(y-14)-2(y-14)=0
(y-14)(y-2)=0
y = 2, 14
...
m.ananth@hotmail.ca